OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, find h[j(2)]

OpenStudy (mathteacher1729):

What is confusing you about this problem?

OpenStudy (anonymous):

j(2)=-2(2)=-4 h(j(2)=h(-4)====? i hope you can find it now

OpenStudy (anonymous):

j(2)= -2(2)= -4 h[j(2)]=h[-4]=3(-4)-1=-13

OpenStudy (anonymous):

did you get it??